OpenStudy (anonymous):

I need help with two Financial Algebra questions. I always give a medal to the best answer. Please, PLEASE help me out! (:

OpenStudy (anonymous):

First question: Ashley is comparing three investment accounts offering different rates. Option A: APR of 4.95% compounding monthly Option B: APR of 4.85% compounding quarterly Option C: APR of 4.895% compounding daily She would like to earn at least a 5% annual yield. Which account(s) will give Alexandra the yield she wants? Options A and B only Option B only Options A and C only Option C only

OpenStudy (anonymous):

Second question: Part 1: Explain which companies would advertise APR over APY and describe how this would help them attract customers. (4 points) Part 2: Create a unique APR (state how often the rate is compounded) and calculate the corresponding APY. Use a comparison of the two rates to verify your answer to part 1 (I only need help with Part 2!)

OpenStudy (amistre64):

well the relationship between apr and apy is: (1+apr/n)^n = (1+apy)

OpenStudy (anonymous):

Okay, so I could say that the APR is 2.35 and the APY is 4.47?

OpenStudy (anonymous):

2.35% and 4.47% is what I meant

OpenStudy (amistre64):

no, they give you the APR, they give you the number of times it occurs in a year; so calculating the relevant APY is just a matter of plugging into the formula.

OpenStudy (anonymous):

Oh, you're talking about question 1!

OpenStudy (amistre64):

for Option A: APR of 4.95% compounding monthly (1 + .0495/12)^12 = 1+ APY yes, and they want that to be 5 or better

OpenStudy (anonymous):

Okay, give me a second and I'll show you what I got

OpenStudy (anonymous):

2.0506386171982448?

OpenStudy (amistre64):

subtract 1, not add 1 :) .050638... is a little over 5%

OpenStudy (anonymous):

Oh, my bad lol So, option A would be the answer?

OpenStudy (amistre64):

A is valid, that narrows it down to 2 choices, see if B is good or bad

OpenStudy (anonymous):

okay hold on (:

OpenStudy (anonymous):

Option B was too low. This is what I got for Option C: 0.0501643943493422 So, the answer would be Option A and C, correct? (:

OpenStudy (amistre64):

very good :)

OpenStudy (anonymous):

Sweet!

OpenStudy (anonymous):

Thanks for the medal!

OpenStudy (amistre64):

you did good work :) now you should be able to do that part 2 just fine.

OpenStudy (anonymous):

Thanks! So, for part two, I can just pick any percentage for the APR and APY I want to create my "unique" APR?

OpenStudy (amistre64):

pick your favorite APR, like 10% choose your favorite compounding rate: 1,000,000 lol and use them to define a corresponding APY value

OpenStudy (anonymous):

lol Okay! May I ask if you could stay? I just want to make sure I am doing this correctly. I really want to get a 100% on this quiz (:

OpenStudy (amistre64):

stay for more questions? or stay to see if you did part 2 correctly?

OpenStudy (anonymous):

Stay to see if I did part 2 correctly

OpenStudy (amistre64):

i might have time for that; make sure you let me know the APR and compounding rate you pcik so i can verify it

OpenStudy (anonymous):

Okay will do. (: It shouldn't take me too long.

OpenStudy (anonymous):

APR: 16.24% Compounding rate: 12 (1+0.1624/12)^12 = 1.1750502449692981 - 1 = 0.1750502449692981

OpenStudy (amistre64):

yep, looks good. i would round it to 0.1751, or 17.51% for the APY

OpenStudy (anonymous):

Okay. Thank you very much! You were extremely helpful!

OpenStudy (amistre64):

youre welcome, and good luck